OpenStudy (anonymous):

The slope of the line passing through P(1,4) and Q(x,1) is -3/5. What is the value of x?

OpenStudy (anonymous):

slope is y2-y1/x2-x1 so in this case it is: -3/5 = 4-1/1-x 3/-5=3/1-x -5 =1-x x=6

OpenStudy (anonymous):

@asdfvg thank you very much!